OpenStudy (anonymous):

The speed of the current in a river is 6 mph. A ferry operator who works that part of the river is looking to buy a new boat for his business. Every day, his route takes him 22.5 miles against the current and back to his dock, and he needs to make this trip in a total of 9 hours. He has a boat in mind, but he can only test it on a lake where there is no current. How fast must the boat go on the lake in order for it to serve the ferry operator’s needs?

OpenStudy (anonymous):

He needs to go 5 mi/hr on the lake for the boat to work on the river also.

OpenStudy (anonymous):

how did you find that out? i need to show the steps

OpenStudy (anonymous):

Let s = the speed of the boat on the lake. s = \[\frac{ 2*22.5 }{ 9 }\] s = \[\frac{ 45 }{ 9 }\] s = 5.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

No prob!